OpenStudy (anonymous):

Find the vector projection of u onto v. Then write u as a sum of two orthogonal vectors, one of which is proj(sub)v(u) u = <-8, 3>, v = <-6, -2>

OpenStudy (anonymous):

\[proj _{v}u\] ^^

OpenStudy (anonymous):

hint u.v =dot product of vectors and |u| denotes magnitude of vector u u.v=|u|(projection of v on u)

OpenStudy (anonymous):

@matricked I'm still a bit confused. And are those intended to be absolute value bars or magnitude bars?

OpenStudy (anonymous):

\[\bf proj_{\vec{v}}\vec{u}=\frac{|\vec{u}|\cos(\theta)}{ |\vec{v}| }\vec{v}=\frac{ \vec{u} \frac{}{} \vec{v} }{ |\vec{v}|^2 }\vec{v}\]

OpenStudy (anonymous):

@RPguy Can you find the projection?

OpenStudy (phi):

Let's say you have a vector u pointing in some direction and another vector v pointing in a different direction. we can draw this picture